The Diagonal of a Rectangle is 25 cm. long and makes an angle of 20 degrees with one side , what are the dimensions of the Rectangle ?

Answer by TimothyLamb(4379)

---
cosA = adj/hyp
cos(20) = a/25
a = 25*cos(20)
a = 23.492315519647711
---
sinA = opp/hyp
sin(20) = b/25
b = 25*sin(20)
b = 8.550503583141717
---
answer:
a = 23.492315519647711 cm
b = 8.550503583141717 cm
---
check:
d = sqrt( a*a + b*b ) = 25
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 4-equations 4-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php